1.      Consider a small economy described by the following equations by the following equations: Y=10,000 mil Eur G=2,000 mil Eur T=2.500 mil Eur, C=500+0.8(Y-T), I=4.000-250r (r denoted in %). Select any number of options. None of the options may be correct as well of them may be correct.

a)     In this economy, private saving, public saving and national saving equal 3,500, 500, 4,000 respectively

b)     In this economy, private saving, public saving, and national equal 1,000, 500, 1,500 respectively

c)      The equilibrium interest rate is 10% p.a

d)     The equilibrium interest rate is 2% p.a

e)     Now, supposing the same economy to be open and NX =4,500-3000*R, the equilibrium real exchange rate is 1.

f)      Do not answer this question





1
Expert's answer
2021-04-08T09:09:04-0400

b)     In this economy, private saving, public saving, and national equal 1,000, 500, 1,500 respectively.

"c=500+0.8(Y-T)"

"C=500+0.8(10000-2500)=6500"

Private saving=Y-T-C

"10000-2500-6500=1000"

Public saving=T-G

"2500-2000=500"

National saving=private saving + public saving

"1000+500=1500"


c)      The equilibrium interest rate is 10% p .a

"Y=C+I+G"

"10000=6500+4000-250r+2000"

"250r=12500-10000"

"r=10" %